OpenStudy (anonymous):

Please help! Show all steps. Will give medal to best answer! Identify the missing term of the geometric sequence: 6,__,54,...

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

@whpalmer4 is it different for geometric sequences?

OpenStudy (anonymous):

6 * 3 = 18 18 * 3 = 54 so it's 18. Trial and error :))

OpenStudy (anonymous):

i'm not sure but here's my method since it's geometric, their ratio with the unknown term (x) will be equal \(\Large \frac{x}{6} = \frac {54}{x}\) solve for "x" \(x^2=(54)(6)\\x= \sqrt{864}\\x=18\) i got same thing XD

OpenStudy (anonymous):

Thank you both!!

OpenStudy (whpalmer4):

You could find the geometric mean of the two: \[GM = \sqrt{a*b} = \sqrt{6*54} = \sqrt{324} = 18\]